a two digit number is three times the sum of each digit. If 45 is addad to the number, the digit reverses. Find the number.
Answer by mananth(16946)

let the number be xy.
10x+y= 3(x+y)
10x+y=3x+3y
7x-2y=0...................1
..
10x+y+45 = 10y+x
9x-9y=-45
x-y=-5....................2
..
multiply equation 2 by -2 and add to equation 1
-2(x-y)+7x-2y=10+0
-2x+2y+7x-2y=10
5x=10
x=2
..
x-y=-5
2-y=-5
y=7
the number is 27